A person's level of blood glucose and diabetes are closely related. Let x be a random variable measured in milligrams of glucose per deciliter (1/10 of a liter) of blood. Suppose that after a 12-hour fast, the random variable x will have a distribution that is approximately normal with mean μ = 85 and a standard deviation of σ = 26. What is the probability that, for an adult after a 12-hour fast, x is more than 46? Round your answer to the nearest thousandth.

0.467

0.217

0.433

0.033

0.933
Answer #1

Solution :

Given that ,

mean = = 85

standard deviation = = 26

P(x >46 ) = 1 - P(x <46 )

= 1 - P[(x - ) / < (46- 85) /26 ]

= 1 - P(z <-1.5 )

= 1 - 0.066 = 0.9332

PROBABILITY = 0.933

Answer = 0.933